OpenStudy (anonymous):

solve the equation x^2=80x+81

OpenStudy (anonymous):

(x-40)^2-1681=0 x=40+-41

OpenStudy (anonymous):

x^2 - 80x -81 = 0 Use the quadratic formula, also will work

OpenStudy (anonymous):

Three ways to solve a quadratic equations: 1. Factorisation 2. Quadratic formula 3. Completing the square

OpenStudy (anonymous):

x^2 -81x+1x-81=0 x(x-81)+1(x-81)=0 (x+1)(x-81)=0

OpenStudy (anonymous):

bnc262, two kinds of solution have been done, so i think it is fair for you to try the last method :)

OpenStudy (anonymous):

when my teacher did these similar problems she got 2 answers is there a reason why

OpenStudy (anonymous):

We also got two answers if you look above. When you solve a quadratic equation expression which is equal to zero, you're actually trying to find all values of x for which that expression is equal to zero. If you look at a graph plot for the equation you're solving, you'll find that, for quadratics, the curve will cut the x-axis anywhere from zero to two times. There are, then, either zero, one or two (real) solutions for x, depending on the nature of your quadratic.

OpenStudy (anonymous):

yes every quadratic equation have 2 answers

OpenStudy (anonymous):

Sorry wali, have to be careful. A quadratic equation also can have just one answer or no answer.

OpenStudy (anonymous):

o yes it con have

OpenStudy (anonymous):

when you say solve the equation, what do you want to solve for?